# Signal Processing for Haptic Surface Modeling: a Review

### v1

@@ 3.4 About the temperature
Temperature is a fundamental dimension of the haptic space, significantly affecting our exploration of the world and our interaction with it. It is often considered an independent property, because the sensors required to measure it (both from human and machine viewpoints) are different and complementary to those needed for other tactile properties [65].
In physics, temperature is clearly defined as one of the fundamental dimensions in the International System of Units [72]. However, in the perceptual domain, human thermal sensitivity depends on the physical characteristics of the object in contact. Specifically, the perceived temperature depends on the thermal transfer between the object itself and the human skin. This thermal interaction involves complex mechanisms that include the thermal conductivity and heat capacity of the materials involved, as well as the surface geometry and the nature of the contact [73].
When two objects of different materials are touched, the perceived temperature is influenced by the rate of heat transfer. For instance, metals typically feel colder than wood at the same temperature because metals have higher thermal conductivity, allowing heat to be transferred away from the skin more rapidly. This concept, known as thermal diffusivity [74], explains why certain materials feel warmer or cooler upon contact despite being at the same physical temperature. Moreover, the perception of temperature is not only a function of the material properties but also involves cognitive processes. Our brain integrates these thermal cues with other sensory information to form a coherent perception of the material, influencing our ability to identify and discriminate between different objects based on their thermal properties.
Given the independence and uniqueness of the temperature with respect to other haptic properties, we do not delve into the details in the scope of this survey. For further analysis, please refer to the following comprehensive review [73].
4 Haptic Surface Modeling and Processing
@@
Signal Year Method Collected signal Bumpiness Friction Roughness Stiffness Exploration Neural-based
2008 [44] F, P, SF ,V Human
2009 [75] F, P, SF ,V Human
[76] A, F Human
2010 [77] F, P, V Human
[78] A Human
[79] A, F Human
2011 [80] A, F, P Human
2012 [81] A Human
2013 [82] A, F, V Human
[83] F, P Human
2014 [84] A, F, SV Human
2015 [85] A, F, V Machine
2016 [86] A, F, P Human
[87] F, P, SV Human
[88] A, F, V Human
2018 [89] A, F, P Human
2020 [90] A, F, P Human
2022 [91] A, F, P Human
[92] F, P Machine
Human-Computer Haptic 2023 [93] F, P, V Machine
2016 [94] Img Human/Machine
2017 [95] Img Human
[96] Vid Human/Machine
2021 [97] Img Human
[98] Img Machine
[99] Img Machine
Machine Haptic 2023 [100] Img Machine
2018 [101] Img, F, A Machine
2020 [102] Img, F, A Machine
Mixed Haptic [103] Img, V, F Not specified
Table 2: A = Acceleration, F = Force, P = Position, SF = Sliding Force, V = Velocity, SV = Sliding Velocity
Haptic surface modeling aims at mapping a physical object surface into the corresponding physical feature space of choice (i.e., roughness, bumpiness, stiffness and friction). Along the modeling, physical features may require some processing, filtering, and adaptation to be transmitted to an haptic device and eventually rendered in the perception space.
@@
As shown in Fig. 4, existing modeling approaches can be divided in two main categories: parametric approaches [43], where the object surface is modeled as a parametric function and the more recent data-driven approaches [25], where the mapping between a surface and the relevant physical features is learnt from real data samples.
@@ 4.1 Parametric approaches
Parametric approaches involve using physical, mathematical, and mechanical principles to simulate the surfaces properties. Due to their simplicity, explicit models were popular in the early studies about haptic rendering [43, 104, 105]. The parameters of the modeling functions can be set through measurements taken from real objects. A glance into these methods can be found in [43]. Notwithstanding their attractive features, parametric models show two main limitations. Firstly, they require some a-priori knowledge of the physical phenomenon to be simulated, and this is not always available. Secondly, except for some simple cases, they require considerable sophistication to make them suitable for real-time rendering, where complex and realistic scenes with non-linear behaviors lead to heavy computation [106]. Due to these problems, such methods have found limited application up to now, and researchers have moved towards data-driven approaches, which do not require any a-priori knowledge of the underlying physical world.
@@ 4.2 Data-driven approaches
The general pipeline of data-driven approaches is to collect the signal generated by a sensor that explores the surface, and map it into a physical counterpart in the feature space. The mapping function is directly learnt from real data samples belonging to a given category. Notably, the learned mapping should be successfully applied to model the haptic features of unseen objects, thus demonstrating its generalisation ability. Data-driven approaches, reported in Tab. 2, can be classified as either computer haptic, which refers to human-computer interaction applications, machine haptic, which refers to robot-based applications, or mixed haptic, which mixes data usually employed by computer haptic or machine haptic, depending on the collected signal.
@@
Applications based on the so-called computer haptic aim at modeling the haptic properties of a surface by mimicking the interaction between the surface and the human finger [75]. The usual pipeline involves the use of sensors, which touch the surface of interest and produce one or multiple one-dimensional signals corresponding to the haptic information, as shown in Fig. LABEL:fig:data_collected (a). The most widely used sensors are the inertial ones, such as accelerometers, probes and force sensors.
@@
As shown in the upper part of Tab. 2, various studies have explored different one-dimensional signals translating them into the physical features. In [44], the authors focused on modeling visco-elastic properties of surfaces, particularly silicon-based and rubber materials. They measured position and force applied simultaneously, using interpolation methods to model transient behaviors. In [75], they utilized a Generalized Maxwell Model for elasticity, focusing on visco-elastic bodies and viscous fluids. In [76], they examined bumpiness and roughness using acceleration and force data collected with a WhAT probe, applying an IIR filter for their modeling. The work in [77] assessed human force sensing capacity and evaluated their previous modeling method [44] by testing on two silicon cylinders. The authors used both static and dynamic models based on radial basis function interpolation. In [78], the authors reduced a 3D high-frequency acceleration signal to a 1D signal using their DFT321 transformation. This method approximates the 3D signal by aligning the module and phase through Discrete Time Fourier Transformation. In [79], they used a sensor fusion technique combining haptic data from a probe with visual data from a camera. They applied the Hertzian contact model to describe the stiffness of objects, achieving a 3D profile from the collected acceleration and force data. Building on [78], [80] used Linear Predictive Coding (LPC) to model time-domain patterns from 3D data reduced to 1D, creating a look-up table for data prediction. Subsequently, in [81], they enhanced the method from [80] using an Auto-Regressive Moving Average (ARMA) model, combining past outputs and weighted moving averages of past inputs to associate time-domain patterns with spatial domains. In [82], data were collected with users moving freely along surfaces. The authors used an Auto-Regressive model, assuming signal stationarity and segmenting signals into chunks using an Auto-PARM algorithm. The study by [83] relied on a PHANToM Premium 1.5/3DOF device and an ATI Nano 17 force/torque sensor to measure force and position, focusing on friction and stiffness. In[84], they continued with DFT321, using an AR model and Auto-PARM algorithm for signal segmentation, and employed Delaunay triangulation for interpolation. The authors of [85] modeled isotropic haptic textures with a frequency-decomposed neural network, demonstrating its applicability to anisotropic textures. The data were collected using a Texanner and processed with a neural network model. In [86], they segmented the data using a bottom-up algorithm and interpolated it with a Radial Basis Function Network (RBFN) to estimate Line Spectrum Frequency coefficients. They analyzed homogeneous anisotropic grain textures, scanning in two directions. In [87], the authors used a force sensor and PHANToM device, developing an interpolation model based on the proxy position of an imaginary contact point. They identified the sliding yield surface through automated palpation. The authors in [88] incorporated Dahl friction model for friction, labeled stiffness with the speed of the tooltip, and used an AR model for roughness. They rendered virtual surfaces with a haptic interface augmented with a Tactile Labs Haptuator. In [89] they followed the segmentation and interpolation methods from their previous work, analyzing bumpiness, friction, roughness, and stiffness using an accelerometer, force sensor, and PHANToM device. Next, the same authors, in [90], proposed another method to collect and process data. They demonstrated that by combining their segmentation framework and a user interface, which guides the data collection process to create a human-in-loop system, the approximation quality of the model increases. In [91], they trained a neural network with contact acceleration data collected through a manual scanning stylus. They used attention-aware 1D CNNs and encoder-decoder networks with Bi-LSTM to capture spatial and temporal dynamics. The authors from [92] used a force-feedback device and load cell to measure force and position, employing Radial Basis Function and Random Forest models for interpolation. Lastly, in [93] they utilized a CatBoost approach, which is a variant of gradient boosting. Here they train decision trees to learn the required mapping function for modeling the objects. A notable mention should be made of the work presented in [78, 80], which introduced the DFT321 method. This technique substantially contributed to the modeling of haptic properties and led to the creation of the first dataset containing data on surface haptic properties [107]. However, the DFT321 approach relies on lookup tables to model surfaces, which limits its ability to generalize effectively to new, unseen surfaces. In contrast, machine learning techniques offer enhanced generalization capabilities, enabling more robust and versatile modeling of haptic properties across a broader range of surfaces.
@@
Concluding, human-computer haptic signals have long been the focus of Haptic Surface Modeling thanks to their simple processing techniques, which allows to immediately characterise the features of a specific point on the surface. However, the same sensed signals (e.g., acceleration, force) have been modeled into different physical features depending on the design of each sensor and the dataset being used. This makes it difficult for researchers to perform comparative studies. As an example, relating how different sensors characterise the physical features of the same surface would help in comparing the different modeling techniques. The other main limitation of human-computer approaches is the difficult in geometrically relating different points in space.
@@
While human-computer haptic applications have been developed to work with humans using one-dimensional signals, machine haptics aims at making a robotic hand deal with objects to accomplish different tasks, such as object manipulation [38]. To achieve this goal, object geometry must be made available. Therefore, machine haptic generally relies on vision-based sensors to obtain 2D high-resolution information in the form of RGB images and haptic/tactile maps, as shown in Fig. LABEL:fig:data_collected (b). The sensors capturing these data rely on a RGB camera recording the deformation of a membrane upon its contact with a surface. The images of the deformed membrane can be directly used as an haptic map or they can be used to estimate the depth of an object representing its shape. In [94], a method for estimating object hardness using a touch sensor and linear regression is proposed, introducing the GelSight sensor. The development and application of the GelSight sensor are thoroughly reviewed in [95]. This paper covers the sensor’s optical system design and algorithms for deriving the haptic physical features from the images. GelSight’s performance in geometry and force measurement is experimentally evaluated, demonstrating its capabilities in capturing friction, roughness, and bumpiness. In [96], a neural network is employed to map an image sequence to a scalar hardness value. This approach measures the hardness of objects with different shapes under loosely controlled contact conditions, utilizing both manual and robotic interactions. In [97] the authors introduce the GelStereo sensor and the method that converts tactile information into geometric deformation of a silicon gel layer, captured using visual sensing methods. The study employs a self-supervised GS-DepthNet estimation technique to enhance the accuracy and resolution of the contact geometry, focusing on roughness and bumpiness. The DelTact sensor, discussed in [98], uses a Gaussian density regression model to estimate roughness and bumpiness from an image. The model solves a point cloud that aligns with prior depth information, ensuring depth consistency and generalizing across sensors with minimal parameter tuning. In [99], the DenseTact model estimates depth images from RGB input images using a DenseNet. This method aims to capture roughness and bumpiness by converting tactile information into depth images. The improved DenseTact2.0 sensor is presented in [100], which includes an encoder-decoder network for shape estimation and force estimation models using transfer learning.
@@
Human-computer and machine haptic applications have traditionally been studied separately due to their different requirements, each facing limitations such as difficult in scaling to the large variety of real materials or inadequate modeling of skin-material interactions. Recently, researchers have proposed methods that integrate the two approaches to model the haptic properties of surfaces. These methods use vision-based sensors or standard cameras to estimate 2D geometric data and inertial sensors for stiffness and friction. In [101], the authors combined data-driven and parametric modeling to estimate the haptic properties of surfaces. They used a photometric stereo algorithm to estimate the bidirectional reflectance distribution function (BRDF), modeling surface texture through data-driven methods for roughness and bumpiness, and physical-based approaches for stiffness and friction. In [102], the model is enhanced by incorporating an LPC-based method to capture vibrations. By merging force and vibration feedback models based on their spectral characteristics, a comprehensive hybrid framework is devised, which allows for more accurate 3D surface reconstruction. In [103], the authors enhanced the data collected in [107] by employing a GelSight sensor to refine the samples of some materials with additional geometric information. Their architecture encodes the GelSight images into texture representation vectors, which are then combined with encoded action representations derived from the user’s force and velocity. This combined representation is used by an acceleration predictor module to predict the desired DFT (Discrete Fourier Transform).

### v2

@@ 3.4 About the temperature
Temperature is a fundamental dimension of the haptic space, significantly affecting our exploration of the world and our interaction with it [chen2009exploring, tiest2010tactual, okamoto2012psychophysical, nellis2008heat]. Temperature works differently from other haptic properties, because the sensors required to measure it (both from human and machine viewpoints) are different and complementary to those needed for other tactile properties [okamoto2012psychophysical].
In physics, temperature is clearly defined as one of the fundamental dimensions in the International System of Units [international2001international]. However, in the perceptual domain, human thermal sensitivity depends on the physical characteristics of the object in contact. Specifically, the perceived temperature depends on the thermal transfer between the object itself and the human skin. This thermal interaction involves complex mechanisms that include the thermal conductivity and heat capacity of the materials involved, as well as the surface geometry and the nature of the contact [ho2018material].
When modeling physical haptic properties, temperature plays a unique role on the interconnection of the haptic properties, but its impact varies for each material category [grimvall1999thermophysical]. For instance, certain types of plastic or organic materials undergo significant changes in haptic properties with changes in temperature. In contrast, materials such as metal, rock, wood, ceramics and fabric remain largely unaffected within temperature ranges that humans can tolerate.
When two objects of different materials are touched, the perceived temperature is influenced by the rate of heat transfer. For instance, metals typically feel colder than wood at the same temperature because metals have higher thermal conductivity, allowing heat to be transferred away from the skin more rapidly. This concept, known as thermal diffusivity [salazar2003thermal], explains why certain materials feel warmer or cooler upon contact despite being at the same physical temperature. Moreover, the perception of temperature is not only a function of the material properties but also involves cognitive processes. Our brain integrates these thermal cues with other sensory information to form a coherent perception of the material, influencing our ability to identify and discriminate between different objects based on their thermal properties. In [chen2009exploring, tiest2010tactual, okamoto2012psychophysical, cai2020thermairglove], a deeper look into the relationship between human haptic perception and the temperature can be found.
Given the uniqueness of the temperature with respect to other haptic properties in the context of surface modeling, we do not delve into the details in the scope of this survey. For further analysis about the temperature related to the whole haptic pipeline, please refer to the following comprehensive review [ho2018material].
4 Haptic Surface Modeling
@@
Approach Year Method Collected signal Bumpiness Friction Roughness Stiffness Exploration Neural-based
2008 [hover2008data] F, P, SF ,V Human
2009 [hover2009data] F, P, SF ,V Human
[guruswamy2009modelling] A, F Human
2010 [hover2010user] F, P, V Human
[landin2010dimensional] A Human
[lang2010measurement] A, F Human
2011 [romano2011creating] A, F, P Human
2012 [culbertson2012refined] A Human
[yim2012shape] F, P Human
2013 [culbertson2013generating] A, F, V Human
[sianov2013data] F, P Human
2014 [culbertson2014modeling] A, F, SV Human
2015 [shin2015data] A, F, V Machine
2016 [abdulali2016data] A, F, P Human
[yim2016data] F, P, SV Human
[culbertson2016importance] A, F, V Human
2018 [abdulali2018data] A, F, P Human
2020 [abdulali2020visually] A, F, P Human
2022 [joolee2022deep] A, F, P Human
[cha2022data] F, P Machine
Human-Computer Haptic
2023 [kumar2023catboost] F, P, V Machine
2016 [yuan2016estimating] Img Human/Machine
2017 [yuan2017gelsight] Img Human
[yuan2017shape] Vid Human/Machine
2021 [cui2021self] Img Human
[du20223d] Img Machine
[do2022densetact] Img Machine
Machine Haptic
2023 [do2023densetact] Img Machine
2018 [shin2018geometry] Img, F, A Machine
2020 [shin2020hybrid] Img, F, A Machine
Mixed Haptic
[heravi2020learning] Img, V, F Not specified
Table 2: We provide a classification of data-driven approaches to haptic surface modeling, highlighting the collected signals, the physical features measured, the exploration method and the neural or deterministic nature of the algorithm used to map the collected signals to the target physical features. A = Acceleration, F = Force, P = Position, SF = Sliding Force, V = Velocity, SV = Sliding Velocity, Img = Image, Vid = Video
Haptic surface modeling consists of mapping the physical properties that characterize a surface onto the corresponding physical feature space of choice (i.e., roughness, bumpiness, stiffness and friction). In this context, properties like force and shape are not directly relevant, as they serve to model perceptual features. Along the modeling, physical features may require some processing, filtering, and adaptation to be transmitted to an haptic device and eventually rendered in the perception space.
@@
As shown in Fig. 3, existing modeling approaches can be divided in two main categories: parametric approaches [basdogan2002haptic], where the object surface is modeled as a parametric function and the more recent data-driven approaches [costes2020towards], where the mapping between a surface and the relevant physical features is learnt from real data samples.
@@ 4.1 Parametric approaches
Parametric approaches involve using physical, mathematical, and mechanical principles to simulate the surfaces properties. Due to their simplicity, explicit models were popular in the early studies about haptic rendering [basdogan2002haptic, pacchierotti2015cutaneous, okamura2001reality][saga2013simultaneous]. The parameters of the modeling functions can be set through measurements taken from real objects. A glance into these methods can be found in [basdogan2002haptic]. Notwithstanding their attractive features, parametric models show two main limitations. Firstly, they require some a-priori knowledge of the physical phenomenon to be simulated, and this is not always available. Secondly, except for some simple cases, they require considerable sophistication to make them suitable for real-time rendering, where complex and realistic scenes with non-linear behaviors lead to heavy computation [peterlik2007algorithm]. Due to these problems, such methods have found limited application up to now, and researchers have moved towards data-driven approaches, which do not require any a-priori knowledge of the underlying physical world.
@@ 4.2 Data-driven approaches
The general pipeline of data-driven approaches is to collect the signal generated by a sensor that explores the surface, and map it into a physical counterpart in the feature space. The mapping function is directly learnt from real data samples belonging to a given category. Notably, the learned mapping should be successfully applied to model the haptic features of unseen objects, thus demonstrating its generalization ability.
We adopt a data-driven approach to modeling, which needs specifying how the data are collected — or rather, explored. The method of exploration affects the meaning and characteristics of the collected data. On the one hand, human exploration involves interacting with the object by applying forces and manipulating it based on pre-existing knowledge of its material and consistency. This type of exploration provides insight into how a final user might perceive and interact with the object. On the other hand, machine exploration applies consistent forces and exploration modalities, irrespective of the object being explored. This results in a more objective and standardized measure of the object’s properties. Thus, human exploration is more suited to capturing subjective perceptions, while machine exploration offers a uniform, objective assessment of the object.
Another key aspect of data-driven approaches is the algorithm that allow to model the collected signals (e.g. force, position, acceleration) into the physical features of a given surface. These algorithms can be either deterministic or neural-based. Deterministic algorithms usually rely on combination of filters like the Auto-Regressive Moving Average (ARMA) to combine multiple signals into a polynomial to model physical features. Neural-based algorithms have employed classic deep network designs like auto-encoders, Convolutional Neural Networks (CNN) to map input (mostly vision-based like images and videos) to physical features. While deterministic approaches provide more control over the process and enhanced explainability, neural-base approaches have shown superior abilities in capturing the relations between the collected signals and physical features.
Data-driven approaches, reported in Tab. 2, can be classified as either computer haptic, which refers to human-computer interaction applications, machine haptic, which refers to robot-based applications, or mixed haptic, which mixes data usually employed by computer haptic or machine haptic, depending on the collected signal.
@@
Applications based on the so-called computer haptic aim at modeling the haptic properties of a surface by mimicking the interaction between the surface and the human finger [hover2009data]. The usual pipeline involves the use of sensors, which touch the surface of interest and produce one or multiple one-dimensional signals corresponding to the haptic information, as shown in Fig. 5(a) (a). The most widely used sensors are the inertial ones, such as accelerometers, probes and force sensors.
@@
As shown in the upper part of Tab. 2, various studies have explored different one-dimensional signals translating them into the physical features. In [hover2008data], the authors focused on modeling visco-elastic properties of surfaces, particularly silicon-based and rubber materials. They measured position and force applied simultaneously, using interpolation methods to model transient behaviors. In [hover2009data], they utilized a Generalized Maxwell Model for elasticity, focusing on visco-elastic bodies and viscous fluids. In [guruswamy2009modelling], they examined bumpiness and roughness using acceleration and force data collected with a WhAT probe, applying an IIR filter for their modeling. The work in [hover2010user] assessed human force sensing capacity and evaluated their previous modeling method [hover2008data] by testing on two silicon cylinders. The authors used both static and dynamic models based on radial basis function interpolation. In [landin2010dimensional], the authors reduced a 3D high-frequency acceleration signal to a 1D signal using their DFT321 transformation. This method approximates the 3D signal by aligning the module and phase through Discrete Time Fourier Transformation. In [lang2010measurement], they used a sensor fusion technique combining haptic data from a probe with visual data from a camera. They applied the Hertzian contact model to describe the stiffness of objects, achieving a 3D profile from the collected acceleration and force data. Building on [landin2010dimensional], [romano2011creating] used Linear Predictive Coding (LPC) to model time-domain patterns from 3D data reduced to 1D, creating a look-up table for data prediction. Subsequently, in [culbertson2012refined], they enhanced the method from [romano2011creating] using an Auto-Regressive Moving Average (ARMA) model, combining past outputs and weighted moving averages of past inputs to associate time-domain patterns with spatial domains. The study by [yim2012shape] relied on a PHANToM Premium 1.5 High Force device and an ATI Nano 17 force/torque sensor to model the shape variation (stiffness) of soft objects by leveraging force and position collected and using a sum of square errors (SSE) fitting function. A second approach employing a similar hardware, presented by [sianov2013data], measures force and position to model friction and stiffness. In [culbertson2013generating], data were collected with users moving freely along surfaces. The authors used an Auto-Regressive model, assuming signal stationarity and segmenting signals into chunks using an Auto-PARM algorithm. In [culbertson2014modeling], they continued with DFT321, using an AR model and Auto-PARM algorithm for signal segmentation, and employed Delaunay triangulation for interpolation. The authors of [shin2015data] modeled isotropic haptic textures with a frequency-decomposed neural network, demonstrating its applicability to anisotropic textures. The data were collected using a Texanner and processed with a neural network model. In [abdulali2016data], they segmented the data using a bottom-up algorithm and interpolated it with a Radial Basis Function Network (RBFN) to estimate Line Spectrum Frequency coefficients. They analyzed homogeneous anisotropic grain textures, scanning in two directions. In [yim2016data], the authors used a force sensor and PHANToM device, developing an interpolation model based on the proxy position of an imaginary contact point. They identified the sliding yield surface through automated palpation. The authors in [culbertson2016importance] incorporated Dahl friction model for friction, labeled stiffness with the speed of the tooltip, and used an AR model for roughness. They rendered virtual surfaces with a haptic interface augmented with a Tactile Labs Haptuator. In [abdulali2018data] they followed the segmentation and interpolation methods from their previous work, analyzing bumpiness, friction, roughness, and stiffness using an accelerometer, force sensor, and PHANToM device. Next, the same authors, in [abdulali2020visually], proposed another method to collect and process data. They demonstrated that by combining their segmentation framework and a user interface, which guides the data collection process to create a human-in-loop system, the approximation quality of the model increases. In [joolee2022deep], they trained a neural network with contact acceleration data collected through a manual scanning stylus. They used attention-aware 1D CNNs and encoder-decoder networks with Bi-LSTM to capture spatial and temporal dynamics. The authors from [cha2022data] used a force-feedback device and load cell to measure force and position, employing Radial Basis Function and Random Forest models for interpolation. Lastly, in [kumar2023catboost] they utilized a CatBoost approach, which is a variant of gradient boosting. Here they train decision trees to learn the required mapping function for modeling the objects. A notable mention should be made of the work presented in [landin2010dimensional, romano2011creating], which introduced the DFT321 method. This technique substantially contributed to the modeling of haptic properties and led to the creation of the first dataset containing data on surface haptic properties [culbertson2014one]. However, the DFT321 approach relies on lookup tables to model surfaces, which limits its ability to generalize effectively to new, unseen surfaces. In contrast, machine learning techniques offer enhanced generalization capabilities, enabling more robust and versatile modeling of haptic properties across a broader range of surfaces.
@@
Concluding, human-computer haptic signals have long been the focus of Haptic Surface Modeling thanks to their simple processing techniques, which allows to immediately characterize the features of a specific point on the surface. However, the same sensed signals (e.g., acceleration, force) have been modeled into different physical features depending on the design of each sensor and the dataset being used. This makes it difficult for researchers to perform comparative studies. As an example, relating how different sensors characterize the physical features of the same surface would help in comparing the different modeling techniques. The other main limitation of human-computer approaches is the difficult in geometrically relating different points in space.
Perception-aware haptic signals
A recently proposed approach in human-computer haptic modeling involves incorporating the user into the modeling process. Unlike previous methods, this set of techniques focuses on modeling haptic properties by taking users’ perceptions into account, relying directly on human judgment to guide the process. One of the earliest works using this approach is presented in [piovarvci2016interaction]. The authors aim to define a model that maps physical stiffness, characterized by surface displacement under applied pressure, to the stiffness perceived by users. To achieve this, they first established a perceptual space that groups similar stiffness-related stimuli. Then, they employed a non-linear model, originally introduced in [pressman2006perception], augmented with terms that account for pressing force to identify the relation between the physical stiffness and its perceived counterpart. In a subsequent study presented in [piovarvci2018perception], the authors constructed a perceptual space for friction and applied a non-metric Multi-Dimensional Scaling (MDS) algorithm, as defined in [wills2009toward], to establish a correlation between the physical property and the perceptual space. Differently from previous method, the authors in [tymms2018quantitative] defined a perceptual space for roughness and bumpiness and create a data-driven model, based on height maps, to map the physical stimuli to the perceptual dimensions. In [degraen2021capturing], the authors sampled 15 different tissues using GelSight sensors and synthetically replicated their characteristics through 3D printing. The study aimed to validate the modeling process by assessing the perceived similarity between the real tissues and their synthetic counterparts. Similarly, in [kurita2023generation], the authors developed a data-driven model for roughness and bumpiness, directly guided by user input. This approach entailed adjusting the model’s input according to user evaluations. In contrast to the previous approaches, other methods aim to model the haptic physical properties indirectly. In [verschoor2020tactile], the authors propose a simulation framework where an hand skin deformation model, guided from a data-driven method, is employed to identify the haptic stimuli and render them to the user. In [shao2023haptic] proposed an indirect method for modeling the haptic properties of surfaces by designing a cognitive model that characterizes how humans interact with physical surfaces.
Involving the user in the loop to model haptic properties is a promising approach for understanding how haptic stimuli can be replicated. However, these methods are challenging to classify and compare with previous approaches, as they do not rely on specific signals to model the properties. Instead, they operate within a psychophysical dimension, requiring distinct evaluation methods that fall outside the scope of this work.
@@
While human-computer haptic applications have been developed to work with humans using one-dimensional signals, machine haptics aims at making a robotic hand deal with objects to accomplish different tasks, such as object manipulation [luo2017robotic]. To achieve this goal, object geometry must be made available. Therefore, machine haptic generally relies on vision-based sensors to obtain 2D high-resolution information in the form of RGB images and haptic/tactile maps, as shown in Fig. 5(a) (b). While these maps can be seen as representations of monodimensional signals, they represent spatially related 1D signals, offering a multidimensional perspective for exploring the collected haptic data. The sensors capturing these data rely on a RGB camera recording the deformation of a membrane upon its contact with a surface. The images of the deformed membrane can be directly used as an haptic map or they can be used to estimate the depth of an object representing its shape. In [yuan2016estimating], a method for estimating object hardness using a touch sensor and linear regression is proposed, introducing the GelSight sensor. The development and application of the GelSight sensor are thoroughly reviewed in [yuan2017gelsight]. This paper covers the sensor’s optical system design and algorithms for deriving the haptic physical features from the images. GelSight’s performance in geometry and force measurement is experimentally evaluated, demonstrating its capabilities in capturing friction, roughness, and bumpiness. In [yuan2017shape], a neural network is employed to map an image sequence to a scalar hardness value. This approach measures the hardness of objects with different shapes under loosely controlled contact conditions, utilizing both manual and robotic interactions. In [cui2021self] the authors introduce the GelStereo sensor and the method that converts tactile information into geometric deformation of a silicon gel layer, captured using visual sensing methods. The study employs a self-supervised GS-DepthNet estimation technique to enhance the accuracy and resolution of the contact geometry, focusing on roughness and bumpiness. The DelTact sensor, discussed in [du20223d], uses a Gaussian density regression model to estimate roughness and bumpiness from an image. The model solves a point cloud that aligns with prior depth information, ensuring depth consistency and generalizing across sensors with minimal parameter tuning. In [do2022densetact], the DenseTact model estimates depth images from RGB input images using a DenseNet. This method aims to capture roughness and bumpiness by converting tactile information into depth images. The improved DenseTact2.0 sensor is presented in [do2023densetact], which includes an encoder-decoder network for shape estimation and force estimation models using transfer learning.
@@
(a) Data-driven approaches require the collection of haptic signals, which can be either (a) mono-dimensional (taken from [culbertson2012refined]) or (b) vision-based (taken from [yang2022touch]). Mono-dimensional signals usually describe a single feature acquired by a sensor, like speed, force, and acceleration. Vision-based signals provide spatially-related 2D high-resolution information of a surface.
@@
Human-computer and machine haptic applications have traditionally been studied separately due to their different requirements, each facing limitations such as difficult in scaling to the large variety of real materials or inadequate modeling of skin-material interactions. Recently, researchers have proposed methods that integrate the two approaches to model the haptic properties of surfaces. These methods use vision-based sensors or standard cameras to estimate 2D geometric data and inertial sensors for stiffness and friction. In [shin2018geometry], the authors combined data-driven and parametric modeling to estimate the haptic properties of surfaces. They used a photometric stereo algorithm to estimate the bidirectional reflectance distribution function (BRDF), modeling surface texture through data-driven methods for roughness and bumpiness, and physical-based approaches for stiffness and friction. In [shin2020hybrid], the model is enhanced by incorporating an LPC-based method to capture vibrations. By merging force and vibration feedback models based on their spectral characteristics, a comprehensive hybrid framework is devised, which allows for more accurate 3D surface reconstruction. In [heravi2020learning], the authors enhanced the data collected in [culbertson2014one] by employing a GelSight sensor to refine the samples of some materials with additional geometric information. Their architecture encodes the GelSight images into texture representation vectors, which are then combined with encoded action representations derived from the user’s force and velocity. This combined representation is used by an acceleration predictor module to predict the desired DFT (Discrete Fourier Transform).